1. (9) Find the general solution to the differential equation. 1) y - 6y +9y = 0 2) y - y - 2y = 0 3) y - 4y + 7y = 0
Answer #1

= Aemx as In each of the cases we consider y = (A+0) a trial solution of the equation and then putting it into the equation wThe given equation is y - by+gy A m² e mx 6 Ame, Mх +9A emx T Aemx (m? - 6m +9 ) =0 m² -6m+g = 0 [ A 70 to] emx &0 m²_ 3m -Hence, the general solution is Ge-x.+62 2x TO where, a and G are arbitrary constants 3 ) is The given equation y - 4y + 7y=0